OpenStudy (anonymous):

Let n e N. Explain what is meant by saying that Pi is a permutation of n = {1, 2, ..., n}. Now let n = 6. How many permutations are there of 6? The permutation Pi of 6 is given below in two line notation. Write Pi in disjoint cycle notation. 1 2 3 4 5 6 1 5 6 3 2 4 How many permutations of 6 are there of the form (a b c) (d e f) when written in disjoint cycle notation? got Pi written in disjoint cycle notation as (2 5)(3 6 4), but can't do the other parts.

OpenStudy (anonymous):

first answer is 6!

OpenStudy (anonymous):

how? please explain

OpenStudy (anonymous):

ok. you have 6 elements to permute. call them {1,2,3,4,5,6}

OpenStudy (anonymous):

you have 6 choices of where to send 1. then you have 5 choices of where to send 2, for a total of 6*5 choices. then 4 choices of where to send 3 etc. so all together you have 6*5*4*3*2*1 choices

OpenStudy (anonymous):

Ok so the answer is 720? thanks. do you know how to do other parts?

OpenStudy (anonymous):

i am thinking of a nice theorem that gives it, but it is not coming to me. so maybe we can just count.

OpenStudy (anonymous):

for the last part?

OpenStudy (anonymous):

i am going to say 80 but i could be wrong.

OpenStudy (anonymous):

sorry, how did you get 80?

OpenStudy (anonymous):

reasoning as follows. for (a,b,c) we have 6 choices for a, then 5 for b and 3 for c but the permutation (abc) is the same as (bca) same as (cab) so we have over counted by a factor of 3, and so for the first cycle there are 6*5*4/3=40 ways. then for the next there are only actually two possibilities. say we are left with {def} then the only three cycles are (def) and (dfe) and 4082=80

OpenStudy (anonymous):

last line should be 40*2=80

OpenStudy (anonymous):

let me think for a moment to see if this is correct.

OpenStudy (anonymous):

4 for C right? and thanks. okay.

OpenStudy (anonymous):

i get the first part, the first cycle but no idea about the second.

OpenStudy (anonymous):

i really want to say there are \[\dbinom{6}{3}\times \dbinom{3}{2}=20\times 3=60\] for C.

OpenStudy (anonymous):

okay i get the first part, but why is it 3C2? answer is 60 not 80?

OpenStudy (anonymous):

because i am wrong. answer to C is 80. 80 different 3 cycles in S6 http://www.math.unl.edu/~bharbourne1/M417Spr03/M417Hmwk7Sols.html

OpenStudy (anonymous):

thanks. i understand it until 5(4)(2)(1) = 40. where im confused as to where the numbers come from. i get that they are b,c,e,f but dont understand why we use those numbers.